# An agent-based model of behaviour change calibrated to reversal learning data

### v1

@@ II.2 Reversal learning experiments
We examined the agent’s behaviour using a simulation of single and serial reversal learning experiments. We followed the experimental setup in Buechel2018 ; Boussard2020 and used their empirical data to calibrate and validate our model. In the single reversal learning experiment on guppies in Buechel2018 , there were two trial periods with 30 trials in the first trial period and 66 trials in the second trial period. Here, we set the second trial period to 30 for uniformity as done similarly in Boussard2020 . We note that we used their empirical data on the large-brained guppies only to make a 1-to-1 comparison with the results they presented in the paper. At every trial, the agent (or the guppy) was given two discs to choose from. During the first trial period, one of the discs had a food reward underneath while the other had none, but during the second trial period, the reward associated with each disc was reversed. We simulated this empirical setup by giving the agent in our model two actions to choose from: A and B. For the first trial period, we set the stimulus $s$ to +1 for action A, signifying the presence of the food reward, and 0 for action B, indicating the absence of the food reward. We then switched the stimulus associated with each action after the reversal point to mimic the reward reversal. The model dynamics of one simulation run of this experiment are illustrated in Figure 3. We also simulated a serial reversal learning experiment which extends the single reversal to a series of reversals as conducted in Boussard2020 . The configurations of the serial reversal learning experiment were similar to the single reversal learning but instead of having only two trial periods separated by one reversal, there were 11 trial periods with ten reversals. These agent-based model parameters are summarized in Table 1. Similar to Buechel2018 ; Boussard2020 , we report the success rate or the fraction of agents that successfully chose the rewarded disc at every trial.
@@
Figure 3: Model dynamics of one simulation run of one agent with a memory decay $\tau=5$ and a logistic function steepness $k=0.5$. (a) The agent selects an action $a_{t}$ (either AA or $B$) at every time step $t$. Highlighted in grey are instances when the agent switches actions. (b) The agent initially receives a stimulus $s_{t}=+1$ for action AA and a stimulus $s_{t}=0$ for action $B$ before the reversal point, but this reward association switches after the reversal. (c) The agent has a memory $m_{t}$ indicating the agent’s preference for each action over time. (d) The agent chooses action $A$ with probability $p_{A}$ based on the stimulus $s_{t}$ and memory $m_{t}$.
II.3 Extended model for serial reversal learning
@@
The result of the serial reversal learning experiments in Boussard2020 revealed a notable characteristic in the agents’ behaviour wherein the agents’ success rate at the end of each trial period decreases over reversals. In our initial simulation runs of the serial reversal learning, we found that our simple two-parameter model was inadequate in capturing this observed phenomenon. Hence, we introduced a third parameter $\theta$ that controls the decay per reversal of the steepness $k$ of the agent’s decision function. Specifically, we decreased the value of $k$ by a factor of $\theta\in[0,1]$ after every reversal period $j$:
$k_{j+1}=k_{j}\times(1-\theta).$ (5)
@@
We can interpret this decay as reducing the impact of information on the agent’s decision by decreasing the steepness of its decision function, which means that higher values of utility $u$ are needed to produce the same probability $p_{a}$.
II.4 Calibration: Approximate Bayesian computation
@@
We used approximate Bayesian computation (ABC) to calibrate our model to empirical data. ABC rejection sampling is a method for estimating the model parameters’ posterior distributions using repeated simulations. Initially, the range of values that each parameter can have is assumed to follow some prior distribution. An ensemble of $N$ models with the parameter values sampled from those priors is then run and the summary statistics of the ensemble are compared with those of the empirical data based on a distance metric. The sampled parameter values are accepted if the distance between the two summary statistics is within the prespecified error threshold $\delta$; otherwise, they are rejected. New parameter values are sampled from the prior distributions and the process repeats until $l$ samples are accepted. These accepted samples characterise the posterior distributions of the model parameters which can be utilised to simulate the observed behaviours from the empirical data. Here, we assumed a uniform prior distribution for each parameter, used the success rate over trials as the summary statistics, and chose the mean squared error as the distance metric.
@@
Table 1: Parameters of the agent-based model and the approximate Bayesian computation used for the single and serial reversal learning simulations.
For the serial reversal learning simulations, the calibration was performed in two stages. The first stage of the calibration process is identical to the one we used for the single reversal learning simulations where the posterior distributions of the memory decay $\tau$ and the steepness $k$ are calculated for the first two trial periods (separated by a single reversal). In the second stage, we computed the posterior distribution of $\theta$ that captures the behaviour across multiple reversals, with the $\tau$ and $k$ parameter values fixed at the maximum a posteriori (MAP) parameter estimate obtained from the first stage. We note that this two-stage calibration procedure is not guaranteed to converge to a maximum likelihood posterior over the joint distribution of $k$, $\tau$, and $\theta$, but nevertheless performs well enough to allow adequate reproduction of the observations after calibration. The parameter ranges we used in our ABC methods are listed in Table 1.
@@ III Results
III.1 Single Reversal Learning
@@
The simulations recovered the observed trajectory of the success rate given parameter values drawn from the posterior distributions of the memory decay $\tau$ and steepness $k$ obtained from the ABC as shown in Figure 4. The success rate shows the average success across 100 agents with the same $\tau=9.7$ and $k=0.26$ parameters. At the start of the first trial period, the success rate is only about 50% because the agents are still unaware of the task at hand. The agents eventually learn which option has the reward after about 20 trials wherein the success rate reaches 80%, which is the criterion set in Buechel2018 to indicate that an agent learned to discriminate the rewarded and non-rewarded options. When the reward is reversed after the first trial period, the success rate drops from about 90% to 10%. This drop in performance indicates the effect of memory on the agents’ decision-making process. Despite the change in the stimulus from the environment, the agents’ memory still favours the previously rewarded option. The agents eventually learn the new reward association over time such that the success rate steadily increases back to 80%.
@@
Figure 4: Model calibration to single reversal learning experiments. The trajectory in (a) shows the success rate over an ensemble of 100 agents in a single reversal learning experiment in which the reward reversal occurs after 30 trials. The black trace is the experimental data from Buechel2018 , the grey traces are the simulation results using all accepted ABC samples, and the orange trace is the simulation result using the MAP parameter estimate. The ABC posterior is shown in (b) with the MAP parameter estimate indicated by the orange point.
III.2 Serial Reversal Learning
@@
As shown in Figure 5a, the first calibration stage of the serial reversal learning simulations also recovered the observed success rate in the first of ten reversals from the experimental data from Boussard2020 . The agents’ success rate goes beyond 80% after several trials, exhibits a huge drop right after the reversal point, and then rises to about 90% eventually. With $\theta=0.12$ obtained from the posterior distribution of $\theta$ after the second calibration stage, the simulated data captures the decline in success rate in the succeeding reversals as depicted in Figure 5c. The success rate after the first reversal goes back up to about 90% but only reaches about 70% after the tenth reversal.
@@
Figure 5: Two-stage model calibration to serial reversal learning experiments. The trajectories in (a) and (c) show the success rate over an ensemble of 100 agents in the first reversal after the first calibration stage and in all ten reversals of a serial reversal learning experiment after the second calibration stage, respectively. The black traces are the experimental data from Boussard2020 , the grey traces are the simulation results using all accepted ABC samples, and the orange traces are the simulation results using the MAP parameter estimate. The ABC posterior is shown in (b) for the first calibration stage and (d) for the second calibration stage with the MAP parameter estimate indicated by the orange point.
@@ IV Discussion
Representation of individual behaviour change in computational models of real-world systems is essential to understanding and predicting phenomena emerging from collective behaviour. However, many behavioural change models rely on arbitrary assumptions unsupported by quantitative empirical data which may limit the validity of simulation results. Here, we demonstrated a simple model of behaviour change that can simulate actual behaviours observed in single reversal learning experiments. We further showed that the model can be easily extended to simulate more complex behaviours in serial reversal learning experiments with longer time scales.
We developed an agent-based model wherein an individual’s behaviour is characterised by three parameters: (1) the memory decay $\tau$ that indicates the extent to which an individual uses previous experiences to make decisions, (2) the steepness $k$ of the logistic decision function that determines an individual’s certainty on the value of given information, and (3) the factor $\theta$ that specifies the decay of an agents certainty about information value over multiple reward reversals. We calibrated these model parameters using ABC given the empirical data from Buechel2018 and Boussard2020 . As shown in Figures 4 and 5, the simulated behaviours using those identified parameters are well-matched to the ensemble statistics observed in the empirical experiments.
@@
Individual behaviours in reversal learning experiments demonstrate a balance between behavioural flexibility (adaptation in response to changing stimuli) and persistence (the tendency for learned behaviour to persist even when rewards are inconsistent). Such a balance is essential for individuals to survive in a dynamically changing world, so it is not surprising that the observed behaviour in guppies in the single reversal learning experiment examined here is consistent with observations of bees Dyer2014 , birds Rayburn-reeves2011 , rodents Le2023 , monkeys Hassett2017 , and humans Rayburn-reeves2011 ; Xia2021 . The decline in success rate observed in the serial reversal experiments we studied suggests a complex mechanism of information processing in which the perceived value of information appears to decay after many fluctuations in the reward signal. We speculate this may be analogous to behavioural phenomena in humans such as declining adherence to protective behaviours during the latter stages of the COVID-19 pandemic Petherick2021 . Designing computational models to have the mechanistic adequacy to simulate such general and specific behavioural change processes could, for example, help better understand animal foraging behaviour Dyer2014 or enable the design of more effective policy interventions during infectious disease outbreaks Petherick2021 . However, to be useful, such models need to have tractable computational complexity and cost, while also maintaining tangible sets of observable inputs and outputs.
@@
In order to calibrate and validate the model to reproduce such real-world behaviours, there is a need for suitable behavioural data. However, because those data are difficult to collect, many existing models rely on assumptions without any empirical support. Moreover, some models that incorporate data rely only on behavioural surveys Durham2012 ; Fierro2013 ; Zhong2013 ; Bayham2015 ; Karimi2015 , which are difficult to scale to larger populations and often characterise individuals’ behavioural intentions that are not necessarily consistent with their actual behaviours Verelst2016 ; Weston2020 . Behavioural data capturing individuals’ actual actions provides a more accurate representation of real-world behaviours. For example, mobility data such as GPS trajectories from mobile phones can be readily collected under different scenarios and conditions Badr2020 ; Xiong2020 ; Ilin2021 . Other indicators of behaviour such as television viewing duration Springborn2015 or home internet usage Zachreson2021 can help quantify ensembles of discrete individual choices. Developing models of individual decision making that can be calibrated to such ensemble data is important for ensuring that simulated behaviours are tangible and valid in the context of the collective behaviour of interest such as the endogenous response to infectious disease outbreaks.
@@
Beyond reproducing real-world behaviours, ensuring that the model parameters that generated such behaviours are identifiable from empirical data is also crucial when designing behavioural change models. Because models are often used to inform decision making, it is desirable to have full and explicit information on what drives a model’s dynamics to improve its interpretability and reliability. By using ABC to calibrate the model, we can characterise the posterior distributions of the model parameters and demonstrate the uncertainty in the model’s dynamics and predictions. Quantifying this model uncertainty could make the analysis and utilisation of models more robust and effective because it provides information on how to interpret their results and on which instances they should be relied on Edeling2021 .
@@
Previous reviews of behavioural change models have also highlighted that many models are not informed by behavioural change theories which provide foundations for understanding, predicting, and influencing individual behaviours. Despite such theories being used in designing and implementing nationwide public health policies and disaster preparedness and response strategies Michie2011 ; Angus2013 ; West2020 , they are rarely incorporated into models that are also developed to support decision making in those scenarios Weston2020 . We designed behaviour in our agent-based model based broadly on the COM-B model for behaviour change which states that capability (C), opportunity (O), and motivation (M) are the key factors that influence behaviour (B) Michie2011 . In our model, the agent’s decision function and internal memory could represent its capability and motivation to behave in a certain way, while the external stimulus could symbolise the opportunity provided by the environment to make the behaviour possible. Our proposed model is a simple computational representation consistent with the COM-B model which can be extended to account for the many other components in each of those factors.
@@
Our study has several limitations. The first is that we only considered a binary set of actions, which many decision-making processes can be reduced to. However, the model can be extended to incorporate more than two actions using other decision functions like a softmax function Sutton2018 . The second limitation is that we used an aggregate behavioural measure to calibrate the model parameters. While those parameters may not be characteristic of each individual’s behaviour, our approach could be extended to represent ensembles of agents with values of individual-level model parameters drawn from population-level hyperdistributions. In practical terms, the approach presented here can be viewed as a technique for building individual-level models that are consistent with ensemble observations. This approach is consistent with typical observables of real-world systems, which are often quantified in aggregate for subpopulations such as communities or demographic strata Springborn2015 ; Badr2020 ; Xiong2020 ; Ilin2021 .

### v2

@@ II.3 Reversal learning experiments
We examined the individual’s behaviour using a simulation of single and serial reversal learning experiments. We followed the experimental setup in Buechel et al., (2018) and Boussard et al., (2020) and used their empirical data to calibrate and validate our model. In the single reversal learning experiment on guppies in Buechel et al., (2018), there were two trial periods with 30 trials in the first trial period and 66 trials in the second trial period. Here, we set the second trial period to 30 for uniformity as done similarly in Boussard et al., (2020). We note that we used their empirical data on the large-brained guppies only to make a 1-to-1 comparison with the results they presented in the paper. At every trial, the individual (or the guppy) was given two discs to choose from. During the first trial period, one of the discs had a food reward underneath while the other had none, but during the second trial period, the reward associated with each disc was reversed. We simulated this empirical setup by giving the individual in our model two actions to choose from: A and B. For the first trial period, we set the stimulus $s$ to +1 for action A, signifying the presence of the food reward, and 0 for action B, indicating the absence of the food reward. We then switched the stimulus associated with each action after the reversal point to mimic the reward reversal. The model dynamics of one simulation run of this experiment are illustrated in Figure 3. We also simulated a serial reversal learning experiment which extends the single reversal to a series of reversals as conducted in Boussard et al., (2020). The configurations of the serial reversal learning experiment were similar to the single reversal learning but instead of having only two trial periods separated by one reversal, there were 11 trial periods with ten reversals. These microsimulation model parameters are summarized in Table 1. Similar to Buechel et al., (2018) and Boussard et al., (2020), we report the success rate or the fraction of individuals that successfully chose the rewarded disc at every trial. To validate the results, we compare the success rate in the simulations ($N=100$) with the success rate in the empirical single ($N=24$) and serial ($N=48$) reversal learning experiments.
@@
Figure 3: Model dynamics of one simulation run of one individual with a memory decay $\tau=5$ and a logistic function steepness $k=0.5$. (a) The individual selects an action $a_{t}$ (either AA or $B$) at every time step $t$. Highlighted in grey are instances when the individual switches actions. (b) The individual initially receives a stimulus $s_{t}=+1$ for action AA and a stimulus $s_{t}=0$ for action $B$ before the reversal point, but this reward association switches after the reversal. (c) The individual has a memory $m_{t}$ indicating the individual’s preference for each action over time. (d) The individual chooses action $A$ with probability $p_{A}$ based on the stimulus $s_{t}$ and memory $m_{t}$.
II.4 Extended model for serial reversal learning
@@
The result of the serial reversal learning experiments in Boussard et al., (2020) revealed a notable characteristic in the individuals’ behaviour wherein the individuals’ success rate at the end of each trial period decreases over reversals. In our initial simulation runs of the serial reversal learning, we found that our simple two-parameter model was inadequate in capturing this observed phenomenon. Hence, we introduced a third parameter $\theta$ that controls the decay per reversal of the steepness $k$ of the individual’s decision function. Specifically, we decreased the value of $k$ by a factor of $\theta\in[0,1]$ after every reversal period $j$:
$k_{j+1}=k_{j}(1-\theta).$ (6)
@@
We can interpret this decay as reducing the impact of information on the individual’s decision by decreasing the steepness of its decision function, which means that higher values of utility $u$ are needed to produce the same probability $p_{a}$.
II.5 Calibration: Approximate Bayesian computation
@@
Similar to previous works, we used approximate Bayesian computation (ABC) to calibrate our microsimulation model to empirical data Van Der Vaart et al., (2015); Grazzini et al., (2017). ABC rejection sampling is a method for estimating the model parameters’ posterior distributions using repeated simulations. Initially, the range of values that each parameter can have is assumed to follow some prior distribution. An ensemble of $N$ models with the parameter values sampled from those priors is then run and the summary statistics of the ensemble are compared with those of the empirical data based on a distance metric. The sampled parameter values are accepted if the distance between the two summary statistics is within the prespecified error threshold $\delta$; otherwise, they are rejected. New parameter values are sampled from the prior distributions and the process repeats until $l$ samples are accepted. These accepted samples characterise the posterior distributions of the model parameters which can be utilised to simulate the observed behaviours from the empirical data. Here, we assumed a uniform prior distribution for each parameter, used the success rate over trials as the summary statistics, and chose the mean squared error as the distance metric. We set the ensemble count $N$ to 100, the error threshold $\delta$ to 0.01, and the sample count $l$ to 1,000. A summary of our ABC parameters is listed in Table 1.
@@
Table 1: Parameters of the microsimulation model and the approximate Bayesian computation used for the single and serial reversal learning simulations.
For the serial reversal learning simulations, the calibration was performed in two stages. The first stage of the calibration process is identical to the one we used for the single reversal learning simulations where the posterior distributions of the memory decay $\tau$ and the steepness $k$ are calculated for the first two trial periods (separated by a single reversal). In the second stage, we computed the posterior distribution of $\theta$ that captures the behaviour across multiple reversals, with the $\tau$ and $k$ parameter values fixed at the maximum a posteriori (MAP) parameter estimate obtained from the first stage. The MAP parameter estimate specifies the mode of the ABC posterior distributions, representing the most likely value of the parameters given the data. We note that this two-stage calibration procedure is not guaranteed to converge to a maximum likelihood posterior over the joint distribution of $k$, $\tau$, and $\theta$, but nevertheless performs well enough to allow adequate reproduction of the observations after calibration.
@@ III Results
III.1 Synthetic reversal learning
To first demonstrate that the ABC algorithm is able to identify ground-truth parameter values, we performed ABC on a synthetic single reversal learning experiment where $\tau=5$ and $k=0.5$. Using these $\tau$ and $k$ parameter values, we ran 24 simulations (representing the 24 large-brained guppies in the single reversal learning experiment in Buechel et al., (2018)) and calculated their success rate which is shown in a solid black line in Figure 4a. We then executed ABC with a configuration similar to that listed in Table 1 for single reversal learning. As exhibited in Figure 4b, the posterior distribution identified by the ABC encapsulates the true parameter values of $\tau=5$ and $k=0.5$ (marked in black), and the MAP parameter estimate is close to these parameter values where $\tau=5.1$ and $k=0.56$ (marked in orange). The success rate trajectory from this calibrated model also fits the success rate trajectory from the synthetic data well as shown in Figure 4a, with a mean squared error of 0.006. On the other hand, Figure 4 also illustrates that uncalibrated models with $\tau$ and $k$ parameter values outside the ABC posterior distribution show success rate trajectories that do not fit the success rate trajectory from the synthetic data, with mean squared errors ranging from 0.015 to 0.079.
III.2 Single reversal learning
@@
The simulations recovered the observed trajectory of the success rate given parameter values drawn from the posterior distributions of the memory decay $\tau$ and steepness $k$ obtained from the ABC as shown in Figure 4. The success rate shows the average success across 100 individuals with the same $\tau=9.7$ and $k=0.26$ parameters. At the start of the first trial period, the success rate is only about 50% because the individuals are still unaware of the task at hand. The individuals eventually learn which option has the reward after about 20 trials wherein the success rate reaches 80%, which is the criterion set in Buechel et al., (2018) to indicate that an individual learned to discriminate the rewarded and non-rewarded options. When the reward is reversed after the first trial period, the success rate drops from about 90% to 10%. This drop in performance indicates the effect of memory on the individuals’ decision-making process. Despite the change in the stimulus from the environment, the individuals’ memory still favours the previously rewarded option. The individuals eventually learn the new reward association over time such that the success rate steadily increases back to 80%. The mean squared error between the observed and simulated success rate trajectories is 0.007.
Figure 4: Model calibration to a synthetic single reversal learning experiment with $\tau=5$ and $k=0.5$. The trajectories in (a) show the success rate when using the $\tau$ and $k$ parameter values marked in (b), where the black mark indicates the ground-truth parameter values, the orange mark indicates the MAP parameter estimate, and the other coloured marks indicate parameter values outside the ABC posterior distribution.
@@
Figure 5: Model calibration to single reversal learning experiments. The trajectory in (a) shows the success rate over an ensemble of 100 individuals in a single reversal learning experiment in which the reward reversal occurs after 30 trials. The black trace is the experimental data from Buechel et al., (2018), the grey traces are the simulation results using all accepted ABC samples, and the orange trace is the simulation result using the MAP parameter estimate. The ABC posterior is shown in (b) with the MAP parameter estimate indicated by the orange point.
III.3 Serial reversal learning
@@
As shown in Figure 5a, the first calibration stage of the serial reversal learning simulations also recovered the observed success rate in the first of ten reversals from the experimental data from Boussard et al., (2020). The individuals’ success rate goes beyond 80% after several trials, exhibits a huge drop right after the reversal point, and then rises to about 90% eventually. With $\theta=0.12$ obtained from the posterior distribution of $\theta$ after the second calibration stage, the simulated data captures the decline in success rate in the succeeding reversals as depicted in Figure 5c. The success rate after the first reversal goes back up to about 90% but only reaches about 70% after the tenth reversal. The mean squared error between the observed and simulated success rate trajectories is 0.008.
@@
Figure 6: Two-stage model calibration to serial reversal learning experiments. The trajectories in (a) and (c) show the success rate over an ensemble of 100 individuals in the first reversal after the first calibration stage and in all ten reversals of a serial reversal learning experiment after the second calibration stage, respectively. The black traces are the experimental data from Boussard et al., (2020), the grey traces are the simulation results using all accepted ABC samples, and the orange traces are the simulation results using the MAP parameter estimate. The ABC posterior is shown in (b) for the first calibration stage and (d) for the second calibration stage with the MAP parameter estimate indicated by the orange point.
@@ IV Discussion
Representation of individual behaviour change in computational models of real-world systems is essential to understanding and predicting phenomena emerging from collective behaviour. All these models must make assumptions on how behaviour changes, but many rely on assumptions with limited support from quantitative empirical data which may limit the validity of simulation results. Here, we demonstrated a simple model of behaviour change that can simulate actual behaviours observed in single reversal learning experiments. We further showed that the model can be easily extended to simulate more complex behaviours in serial reversal learning experiments with longer time scales.
We developed a microsimulation model wherein an individual’s behaviour is characterised by three parameters: (1) the memory decay $\tau$ that indicates the extent to which an individual uses past experiences to make decisions, (2) the steepness $k$ of the logistic decision function that determines an individual’s certainty on the value of given information, and (3) the factor $\theta$ that specifies the decay of an individual’s certainty about information value over multiple reward reversals. We calibrated these model parameters using ABC given the empirical data from Buechel et al., (2018) and Boussard et al., (2020). As shown in Figures 4 and 5, the simulated behaviours using those identified parameters are well-matched to the ensemble statistics observed in the empirical experiments.
@@
Individual behaviours in reversal learning experiments demonstrate a balance between behavioural flexibility (adaptation in response to changing stimuli) and persistence (the tendency for learned behaviour to persist even when rewards are inconsistent). Such a balance is essential for individuals to survive in a dynamically changing world, so it is not surprising that the observed behaviour in guppies in the single reversal learning experiment examined here is consistent with observations of bees Dyer et al., (2014), birds Rayburn-Reeves et al., (2011), rodents Le et al., (2023), monkeys Hassett and Hampton, (2017), and humans Rayburn-Reeves et al., (2011); Xia et al., (2021). The decline in success rate observed in the serial reversal experiments we studied suggests a complex mechanism of information processing in which the perceived value of information appears to decay after many fluctuations in the reward signal. During the COVID-19 pandemic where non-pharmaceutical interventions were repeatedly imposed, lifted, and reinstated, a similar decline in adherence to costly behaviours such as physical distancing was observed Petherick et al., (2021). Designing computational models to have the mechanistic adequacy to simulate such general and specific behavioural change processes could, for example, help better understand animal foraging behaviour Dyer et al., (2014) or enable the design of more effective policy interventions during infectious disease outbreaks Petherick et al., (2021). However, to be useful, such models need to have tractable computational complexity and cost, while also maintaining tangible sets of observable inputs and outputs.
@@
In order to calibrate and validate the model to reproduce such real-world behaviours, there is a need for suitable behavioural data. However, because those data are difficult to collect, many existing models rely on assumptions with limited empirical support. Models that do incorporate behavioural data rely mainly on behavioural surveys Durham and Casman, (2012); Fierro and Liccardo, (2013); Zhong et al., (2013); Bayham et al., (2015); Karimi et al., (2015), which are difficult to scale to larger populations and often characterise individuals’ behavioural intentions that are not necessarily consistent with their actual behaviours Verelst et al., (2016); Weston et al., (2020). Behavioural data capturing individuals’ actual actions provides a more accurate representation of real-world behaviours. For example, mobility data such as GPS trajectories from mobile phones can be readily collected under different scenarios and conditions Badr et al., (2020); Xiong et al., (2020); Ilin et al., (2021). Other indicators of behaviour such as television viewing duration Springborn et al., (2015) or home internet usage Zachreson et al., (2021) can also help quantify ensembles of discrete individual choices. Developing models of individual decision making that can be calibrated to such ensemble data is important for ensuring that simulated behaviours are tangible and valid in the context of the collective behaviour of interest such as the endogenous response to infectious disease outbreaks.
@@
Beyond reproducing real-world behaviours, ensuring that the model parameters that generated such behaviours are identifiable from empirical data is also crucial when designing behavioural change models. Because models are often used to inform decision making, it is desirable to have full and explicit information on what drives a model’s dynamics to improve its interpretability and reliability. By using ABC to calibrate the model, we can characterise the posterior distributions of the model parameters and demonstrate the uncertainty in the model’s dynamics and predictions. Quantifying this model uncertainty could make the analysis and utilisation of models more robust and effective because it provides information on how to interpret their results and on which instances they should be relied on Edeling et al., (2021).
@@
Previous reviews of behavioural change models have also highlighted that many computational models are not informed by behavioural change theories which provide foundations for understanding, predicting, and influencing individual behaviours. Despite such theories being used in designing and implementing nationwide public health policies and disaster preparedness and response strategies Michie et al., (2011); Angus et al., (2013); West et al., (2020), they are rarely incorporated into models that are also developed to support decision making in those scenarios Weston et al., (2020). We designed behaviour in our microsimulation model based on the COM-B model of behaviour change which states that capability (C), opportunity (O), and motivation (M) are the three key components that influence behaviour (B) Michie et al., (2011). Our proposed microsimulation model provides a simple computational representation of behaviour whose components can be mapped to those of the COM-B model, and which can be modified or extended to account for the multiple other factors described in psychological theories of behaviour change.
@@
Our study has several limitations. The first is that we only considered a binary set of actions, which many decision-making processes can be reduced to. However, the model can be extended to incorporate more than two actions using other decision functions like a softmax function Sutton and Barto, (2018). The second limitation is that we used an aggregate behavioural measure to calibrate the model parameters. While those parameters may not be characteristic of each individual’s behaviour, our approach could be extended to represent ensembles of individuals with values of individual-level model parameters drawn from population-level hyperdistributions. In practical terms, the approach presented here can be viewed as a technique for building individual-level models that are consistent with ensemble observations. This approach is consistent with typical observables of real-world systems, which are often quantified in aggregate for subpopulations such as communities or demographic strata Springborn et al., (2015); Badr et al., (2020); Xiong et al., (2020); Ilin et al., (2021).